myininaya (myininaya):

Cal Question Show that cos(arctan(sin(arccot(x)))=sqrt{frac{x^2+1}{x^2+2}}

myininaya (myininaya):

\[\cos(\arctan(\sin(arccot(x)))=\sqrt{\frac{x^2+1}{x^2+2}} \]

OpenStudy (anonymous):

what are you kidding?

hero (hero):

myininaya, if I had to put money on it, I'd bet you already have this solved

myininaya (myininaya):

lol this is a question in the book i just thought it would fun

myininaya (myininaya):

no i was just looking and seen it and thought i would give it a try

OpenStudy (anonymous):

actually i bet i can do it without any calculus at all yes? you have an odd idea of fun

myininaya (myininaya):

i was thinking that satellite

OpenStudy (anonymous):

this question looks ridiculous!

myininaya (myininaya):

i know

hero (hero):

lol

myininaya (myininaya):

thats why i chose it

OpenStudy (anonymous):

oh i lied

OpenStudy (anonymous):

i thought i could just work from the inside out but i cannot. so i guess you do ... would you like my guess?

hero (hero):

myininaya, just post the solution already.

myininaya (myininaya):

i just started working on it

OpenStudy (zarkon):

actually pretty easy

OpenStudy (anonymous):

my guess is take the derivative of both sides. see that you get the same thing. then evaluate somewhere, see that you get the same thing. that is my guess

OpenStudy (anonymous):

i just got it.

OpenStudy (zarkon):

just make 2 right triangles

OpenStudy (anonymous):

zarkon's right, its not that bad.

OpenStudy (anonymous):

i have them on my paper. but what do you do with.. oh never mind. i am a moron

OpenStudy (anonymous):

last step is \[\cos(\arctan(\frac{1}{\sqrt{x^2+1}})\] if i am not mistaken

OpenStudy (anonymous):

right.

hero (hero):

Can somebody post it on a pdf?

OpenStudy (anonymous):

i can write it out maybe, it is not bad

myininaya (myininaya):

easy

OpenStudy (anonymous):

start with \[\sin(\tan^{-1}(x))=\frac{1}{\sqrt{x^2+1}}\]

OpenStudy (anonymous):

i gotta learn to do that. do you scan these?

myininaya (myininaya):

i'm a scanner

myininaya (myininaya):

with a scanner

OpenStudy (anonymous):

lol

myininaya (myininaya):

lol

OpenStudy (anonymous):

got it. oh look some one posted a nice integral for you to do!

myininaya (myininaya):

i scan the hell out of things

myininaya (myininaya):

zarkon remember it is my favorite thing to do to make right traingles lol

OpenStudy (zarkon):

yep...fun for all ages :)

OpenStudy (anonymous):

go get 'em. i'm off.

myininaya (myininaya):

i could give this to my trig students why would you want to show this using cal?

myininaya (myininaya):

i don't know if that is possible to use cal for this problem

OpenStudy (zarkon):

doing it once is good for trig sub...but having to do it twice..i don't see a point